OpenStudy (tmostiller):

Simplify each expression. Use Positive exponents. (x^-2y^-4x^3)^-2 the answer I put was (x-2y-4x3)=1/(x-2y-4x3)^2 Please tell me what I got wrong. I am not good at Algebra AT ALL

OpenStudy (acespeedfighter):

HI i know you, your cool god loves you god bless you, you thy be blessed ʕ•ᴥ•ʔ

OpenStudy (tmostiller):

:)

OpenStudy (sshayer):

\[\frac{ 1 }{( \frac{ x^3 }{ x^2y^4 } )^2}=\frac{ 1 }{ \frac{ x^2 }{ y^8 } }=\frac{ y^8 }{ x^2 }\]